Ali has 20 gallons of 20% sulfuric acid solution. How much 100% sulfuric acid solution must add to make a solution that is steady 50% sulfuric acid.
Answer by Theo(13342)

let y = the number of gallons of 100% acid solution that you have to add to get a 50% acid solution.

you have 20 gallons of 20% acid solution.

your equation will be:

.2 * 20 + y = .5 * (20 + y)

20 + y is the total amount of gallons of solution that you will have when you are done.

simplify this equation to get:

4 + y = 10 + .5 * y

subtract .5 * y from both sides of this equation and subtract 4 from both sides of this equation to get:

y - .5 * y = 10 - 4

simplify to get:

.5 * y = 6

divide both sides of this equation by .5 to get:

y = 12

you will need 12 gallons of 100% acid solution added to 20 gallons of 20% acid solution to get 32 gallons of 50% acid.

the number of gallons of pure acid in the new solution will be .2 * 20 + 12 which is equal to 4 + 12 = 16

16 / 32 = 50%

numbers look good.

your solution is that you will need to add 12 gallons of 100% sulfuric acid to 20 gallons of 20% sulfuric acid to make 32 gallons of 50% sulfuric acid.